# GPU Cluster Scheduling for Network-Sensitive Deep Learning

### v1

@@ 3. MOTIVATION
A typical DL cluster today comprises a large number of GPUs communicating via different types of network interconnects. These interconnects may be (i) intra-machine like NVSwitch (nvlink-nvswitch, ); (ii) intra-rack connected via Infiniband; or, (iii) inter-rack connected via modern network (Ethernet) switches. Such a hierarchical network structure is depicted in Fig. 1. In this figure, a job mapped to GPUs on the same machine is shown in red; a job running on the same rack (but on different machines) is marked in blue; and a job mapped to GPUs across the network is shown in grey. The communication overhead is lowest for the jobs highlighted in red, followed by those in blue and then grey.
Figure 1. A typical (hierarchical) datacenter n/w.
3.1. Communication overhead in DDL Clusters
@@
DDL training throughput is heavily influenced by communication speed and latency due to the significant volume of data transfer among DNN workers during gradient synchronization. As highlighted in (kungfu, ), an 8-GPU server engaged in training a ResNet50 model produces 4 GB of gradients per second. Under layer-wise model synchronization, starting from the output layers under back-propagation: (i) gradients (of the loss training-objective based on the GPUs’ assigned training-data batches) are computed with respect to parameters of the current layer, (ii) these gradients are then exchanged among the GPUs, and finally, (iii) the current layer’s parameters are simultaneously updated by each GPU. Hence, multiple DDL models can quickly overwhelm the link bandwidth. This underscores the importance of high-speed and low-latency interconnects between GPUs for mitigating communication-related stalls. Previous studies such as (pipedream, ; stash, ) reveal that the communication overheads can be as high as 5×5$\times$5 × the actual compute time. To circumvent such substantial overheads, schedulers like Tiresias strive to consolidate DDL jobs on the most proximal GPUs possible.
Figure 2. Single iteration running time for models consolidated on the same machine, rack, and across the network
3.2. Job consolidation in DL cluster
@@
Cluster schedulers like Tiresias (tiresias, ) aim to reduce communication overhead by consolidating job placement, and assigning GPUs to a job within the fewest machines possible. However, due to multi-tenancy, such consolidation can lead to lengthy queueing delays, depending on the cluster’s load. Individual jobs may reject resource offers generated by the cluster scheduler if the offered GPUs are not located on the same machine or rack, especially when GPU demand is high. As a result, jobs remain queued, increasing their job completion time (JCT).
@@
However, not all jobs benefit equally from consolidation, as network overhead varies across DNN models. More specifically, jobs which are less sensitive to network quality (less ‘network-sensitive’ models) can tolerate less stringent network consolidation (consolidation on physically distant GPUs) compared to highly ‘network-sensitive’ jobs. To illustrate this, in Fig. 2, we simulate the single iteration DDL running times using ASTRA-sim for various models with different network placements. As expected, in general, consolidation on physically proximal GPUs yields lower run times. However, we observe significant variations in communication overheads across network tiers for different DNN models (also shown in Table 1). Thus, model network sensitivities could act as a potentially relevant metric in determining the priority of each job with respect to consolidation: the higher the network sensitivity of a model, the more it should be prioritized to get an allocation with physically proximal (consolidated) GPUs.
@@
In this direction, Tiresias, seeks to categorize network-sensitive jobs by employing model ”skew,” which is defined as the ratio of the largest tensor size to the overall model size. Tiresias consolidates jobs with high skew, while other jobs accept any resource offer they receive. However, using ASTRA-sim to analyze the communication overheads of various models under different network placements and high speed networking hardware reveals weak correlation between skew and network sensitivity as shown in Table 1. For e.g. ResNet18 has low skew but very high network level communication overhead and VGG11 has high skew but low communication overhead. This could be due to the rapid advancements in state-of-the-art (SOTA) network hardware.
@@
VGG11 (vgg, ) High 1% 6% 7%
AlexNet (alexnet, ) High 2% 13% 100%
MobileNetV3 (mobilenet, ) High 42% 940% 19592%
ResNet18 (resnet, ) Low 7% 116% 2749%
ResNet50 (resnet, ) Low 12% 12% 38%
BERT large (bert, ) Low 8% 23% 715%
Table 1. Communication overhead as a % of compute time.
3.3. Improvements in SOTA network hardware
@@
Recent advancements in network hardware have introduced remarkably high network bandwidth with exceptionally low latency. Hardware manufacturers, such as NVIDIA, are now offering network switches that deliver several-fold improvements in QoS compared to previous-generation hardware. For intra-machine networking, NVSwitch (nvlink-nvswitch, ) provides a bandwidth of up to 900 Gb/s, surpassing PCIe’s bandwidth of only up to tens of GB/s combined with high latency (li2019evaluating, ). Furthermore, for Infiniband-based intra-rack networking, NVIDIA quantum switches now offer up to 400 Gb/s bandwidth per port with ultra-low latency using GPU direct RDMA. In contrast, standard Infiniband offers only up to 50 Gb/s bandwidth while experiencing relatively high latency (up to tens of microseconds) (Infiniband-eval, ). Finally, NVIDIA spectrum Ethernet switches provide up to 800 Gb/s bandwidth for inter-rack networking. However, current schedulers lack awareness of such high-speed network hardware, which could be harnessed for improved performance. This underscores the demand for a novel approach to DL cluster scheduling.
3.4. Limitations of current SOTA DL cluster schedulers
@@
Current SOTA schedulers like Tiresias and Gandiva (gandiva, ) have several shortcomings such as: (i) Leveraging only limited knowledge of the data center network: due to the lack of awareness of the various network tiers, for e.g., the medium-bandwidth GPU direct RDMA that exists within a rack, these schedulers often end up making a sub-optimal job placement decision, in this case, mapping the job across multiple racks. (ii) A strict consolidation policy: restricting the job placement to either the fewest machines possible (for high skewed models) may lead to high queueing delays in case the former is not available. This necessitates an intermediate approach, allowing for a gradual and graceful ‘relaxation’ of consolidation, particularly for moderately sensitive jobs based on queueing delay and network sensitivity. (iii) Sub-optimal job priority: Tiresias employs a generalized version of Least Attained Service (LAS) (las, ) called ”Discretized 2D-LAS” (2DAS) (tiresias, ) to determine job priority. The 2DAS value for a job is calculated by multiplying the time for which it has been running by the number of GPUs it utilizes. Tiresias prioritizes jobs based on their 2DAS values, with lower values indicating higher priority. Consequently, a job with the highest 2DAS value is least likely to be scheduled for execution and most likely to be preempted out. However, a job with unfavorable network placement and high network sensitivity may have been running for an extended period but may not have made significant progress due to the excessive network overhead associated with its placement. This would lead to a high 2DAS value and make it less likely to receive resource offers that are consolidated, as jobs with lower 2DAS values would receive the most favorable offers.
@@
3.5. A high-fidelity simulator for DDL scheduling research
@@
The exorbitant cost of deep learning GPUs, with several GPUs required to create a large deep-learning cluster, poses a significant challenge in conducting real-world experiments. This, combined with the need to run such experiments for extended periods, can escalate the cost to the hundreds of thousands of dollars. Even utilizing public cloud resources may prove uneconomical as deep learning machines like the AWS P4 instances incur hourly charges of more than \$32/hour (aws_pricing, ). Further complicating matters, research based on cutting-edge network hardware necessitates a simulation platform due to its limited availability. To the best of our knowledge, no public cloud provider offers access to such hardware. To address these challenges, research efforts like (rashidi2022themis, ) have turned to ASTRA-sim as a viable tool for simulating the latest network hardware. However, ASTRA-sim is a single DDL job simulator.
@@
Existing multi-job DDL cluster scheduler simulators, such as (themis-nsdi, ), suffer from low accuracy due to their unrealistic approach to network overhead modeling. For example, (themis-nsdi, ) implements static slowdown penalties of job placement on the rack and the network irrespective of the DNN model or network hardware. But to achieve high confidence in the simulation results, a DL cluster scheduler simulator must dynamically determine the accurate network overhead of a job based on specific placement within the cluster and the type of network hardware used to accurately reflect real-world conditions. In response to this requirement, we propose a multi-job DL cluster simulator that employs ASTRA-sim to accurately determine network overhead.
4. DALLY DESIGN
@@
We now describe Dally, a novel network-placement sensitive DL cluster scheduler that leverages the principles of delay scheduling and novel network-sensitive preemption policy. In the following sections, we delve into the details of our system design, including the simulation methodology employed to evaluate its effectiveness.
4.1. Overview
4.1.1. System goal
@@
The system processes a stream of DL/DDL jobs that need to be scheduled for execution. These jobs exhibit heterogeneity in terms of the DNN model, the number of iterations required to attain the target accuracy, the compute time per iteration, and the GPU demand per job. The system may encounter jobs arriving in batches so that the aggregate GPU demand across all jobs would normally surpass the total number of GPUs available in the cluster. Alternatively, jobs may arrive according to a Poisson distribution. Our goal is to construct a DL cluster scheduler that aligns with the target objectives and assumptions elaborated upon in the subsequent sections. This scheduler will function within a cluster comprising homogeneous GPUs connected via latest-generation network hardware in a hierarchical structure, as discussed in Section 3.
4.1.2. Objective
@@ 4.1.3. Constraints
  3. (3)
    Unknown job submission pattern: The scheduler lacks information about the exact order and frequency of job arrivals.
  4. (4)
@@
4.1.4. System workloads
We employ the cluster job trace made available by SenseTime (sensetime, ) through the work presented in (lucid-asplos23, ). The traces are more recent than those utilized in (philly_trace_analysis, ) and provide a more realistic representation of a real-world DL cluster with multiple jobs being submitted over an extended period. This trace encompasses individual job arrivals alongside pertinent metadata information for each job.
4.2. Simulation methodology
4.2.1. Simulator design
ArtISt-sim builds upon a prior work, Themis (themis-nsdi, ), which is a DL cluster scheduler simulator, and ASTRA-sim. The choice of Themis is based on its unique capability to simulate DL jobs at an iteration level, i.e., a feature not present in other simulators like (lucid-asplos23, ; chronus-socc21, ; tiresias, ). This granular simulation approach enables us to precisely track ”job progression” for preemption and subsequent placement decisions throughout the job’s lifetime. However, as previously discussed, the network slowdown in the original work is encoded statically. To address this limitation, in ArtISt-sim, we modify Themis to invoke ASTRA-sim for each new job placement (refer Fig. 3).
For executing ASTRA-sim with specific network placements, ArtISt-sim dynamically generates the necessary input – workload, system, and network topology files and passes them to ASTRA-sim (blue arrow in Fig. 3). These generated files delineate the ”portion” of the data-center network utilized by the worker GPUs. For instance, in Fig. 1, the grey colored job occupies 8 GPUs across two machines connected via a network link. To simulate ASTRA-sim for such a placement, ArtISt-sim produces the 8-GPU configuration files with a two-tier hierarchy representing the placement. The configuration files specify that the intra-machine GPUs are connected via NVSwitch in a ring topology, and the two machines are linked via a network (Ethernet) switch using a switch topology. Other ASTRA-sim configurations, such as collectives, are also dynamically generated. ASTRA-sim then accurately simulates the job for a single DDL iteration based on the placement and relays the communication overhead back to ArtISt-sim (blue arrow in Fig. 3), which subsequently utilizes this communication overhead to simulate the job’s progression (iterations) until it is either interrupted due to preemption or successfully completes its execution. Note that the compute time per iteration is extracted from the workload trace, and ASTRA-sim serves solely to calculate the network overhead for the specific placement. Network overhead is calculated by first finding the percentage of exposed communication time to the compute time in the ASTRA-sim simulation. The percentage value is applied to the iteration time from the trace to calculate the network overhead and the actual iteration time in ArtISt-sim.
Figure 3. Simulation design
4.2.2. Simulation calibration and accuracy
To ensure the fidelity of our simulations, we employ a calibration process that incorporates real-world measurements into ArtISt-sim. We execute each model listed in Table 1 on an 8-GPU machine connected via NVSwitch and calculate the associated network overhead. Along with this, we simulate the model with the same network configurations (NVSwitch) in ASTRA-sim. Subsequently, we scale the workload configuration file of ASTRA-sim (for each model) to eliminate the discrepancy between simulated and actual network overhead. These modified workload files are then used for all ASTRA-sim calls. The ”calibrated” simulator is now ready to simulate overhead arising from rack and network-level placements.
For the verification of the various baseline scheduler implementations in ArtISt-sim, we rely on the fact that Themis has already been validated in (themis-nsdi, ). Furthermore, we faithfully reproduce the design principles of the Tiresias and Gandiva schedulers as outlined in their respective publications.
4.2.3. Implementation
We have implemented ArtISt-sim in about 2000 lines of java. [^2] Upon acceptance of this work, we will freely provide comprehensive details of all configurations along with the open-source code. ArtISt-sim requires the network configuration parameters for intra-machine, intra-rack, and inter-rack interconnects. These parameters encompass the network topology, bandwidth, latency, collective communication algorithm, and other relevant details. Additionally, the user needs to specify the number of machines per rack and the total number of racks in the cluster.
4.3. Design of the Dally scheduler
@@
Dally is designed to minimize cloud expenditures in both public and private cloud environments. To achieve this objective, Dally uses delay scheduling for job placement and employs a novel preemption strategy that incorporates network sensitivity. When a job is submitted to the cluster, the global scheduler (refer to Fig. 4) initially places it into a wait queue. The scheduler then periodically generates ”resource offers” to jobs in the wait queue based on the availability of cluster resources. If a job’s local scheduler (Fig. 4) decides to accept the offered resources, it is moved to the run queue, where it executes until it is either preempted or completes its execution. Upon preemption, the job is reintroduced into the wait queue and remains there until it accepts another resource offer. An instance of Dally’s mechanisms governing preemption and job placement is elaborated upon in the following sections.
Figure 4. Scheduling scheme.
4.3.1. Preemption priority

### v2

@@ III MOTIVATION
A typical DL cluster today comprises a large number of GPUs communicating via different types of network interconnects. These interconnects may be (i) intra-machine like NVSwitch [nvlink-nvswitch]; (ii) intra-rack connected via Infiniband; or, (iii) inter-rack connected via modern network (Ethernet) switches. Such a hierarchical network structure is depicted in Fig. 2. In this figure, a job mapped to GPUs on the same machine is shown in red; a job running on the same rack (but on different machines) is marked in blue; and a job mapped to GPUs across the network is shown in grey. The communication latency is lowest for the jobs highlighted in red, followed by those in blue and then grey.
Figure 2: A typical (hierarchical) datacenter network (n/w).
III-A Communication Latency in DDL Clusters
@@
The effectiveness of DDL training throughput is heavily influenced by the essential role of communication speed and latency due to the significant volume of data transfer among DNN workers during gradient synchronization. As highlighted in [kungfu], an 8-GPU server engaged in training a ResNet50 model produces 4 GB of gradients per second. Also, previous studies such as [pipedream, stash] reveal that the communication latency can be as high as $5\times$ the actual compute time. Under a layer-wise model synchronization regime, starting from the output layers under back-propagation: (i) gradients (of the loss training-objective based on the GPUs’ assigned training-data batches) are computed with respect to parameters of the current layer, (ii) these gradients are then exchanged among the GPUs, and finally, (iii) the current layer’s parameters are simultaneously updated by each GPU. Hence, multiple DDL models can quickly overwhelm the link bandwidth. This underscores the importance of high-speed and low-latency interconnects between GPUs for mitigating communication-related stalls.
III-B Job Consolidation in DL Clusters
@@
Cluster schedulers like Tiresias [tiresias] aim to reduce communication latency by consolidating job placement, and assigning GPUs to a job within the fewest machines possible. However, due to multi-tenancy, such consolidation can lead to lengthy queueing delays, depending on the cluster’s load. Individual jobs may reject resource offers generated by the cluster scheduler if the offered GPUs are not located on the same machine or rack, especially when GPU demand is high. As a result, jobs remain queued, increasing their job completion time (JCT).
@@
However, not all jobs benefit equally from consolidation, as network latency varies across DNN models. More specifically, jobs which are less sensitive to network quality (less ‘network-sensitive’ models) can tolerate less stringent network consolidation (consolidation on physically distant GPUs) compared to highly “network-sensitive” jobs. To illustrate this, in Fig. 1, we simulate the single iteration DDL running times using ASTRA-sim for various models with different network placements. Each model is simulated using 8 GPUs wherein the GPUs are placed – (i) within the same machine, (ii) within the same rack, and (iii) distributed across racks. As expected, in general, consolidation on physically proximal GPUs yields lower run times. However, we observe significant variations in communication latency across machine, rack and network for different DNN models. Table I illustrates the communication time (latency) expressed as a percentage of the compute (GPU) time. The significant variations in latency, such as ResNet-18 exhibiting a notable rack level latency, stem from the unique characteristics of individual models or their network sensitivity. Prior work [stash] has examined the features of DNN architectures that influence latency behavior in relation to the underlying network topology. Thus, model network sensitivities could act as a potentially relevant metric in determining the priority of each job with respect to consolidation: the higher the network sensitivity of a model, the more it should be prioritized to get an allocation with physically proximal (consolidated) GPUs.
@@
In this direction, Tiresias [tiresias], seeks to categorize network-sensitive jobs by employing model “skew,” which is defined as the ratio of the largest tensor size to the overall model size. Tiresias consolidates jobs with high skew, while other jobs accept any resource offer they receive. The underlying assumption is that skew accurately indicates network sensitivity. However, using ASTRA-sim to analyze the communication latency of various models under different network placements and high speed networking hardware reveals weak correlation between skew and network sensitivity as shown in Table I. For example, ResNet18 has low skew but very high network level communication latency and VGG11 has high skew but low communication latency. This could be due to the rapid advancements in state-of-the-art (SOTA) network hardware which has changed the various latency incurred during gradient exchange. A plausible relationship between the DNN model and communication latency is that the latter depends on the gradient size per layer and the underlying network characteristics such as bandwidth and latency [stash].
@@
VGG11 [vgg] High 1% 6% 7%
AlexNet [alexnet] High 2% 13% 100%
MobileNetV3 [mobilenet] High 42% 940% 19592%
ResNet18 [resnet] Low 7% 116% 2749%
ResNet50 [resnet] Low 12% 12% 38%
BERT large [bert] Low 8% 23% 715%
TABLE I: Communication latency as a % of compute time.
III-C Improvements in SOTA Network Hardware
@@
Recent advancements in network hardware have introduced remarkably high network bandwidth with exceptionally low latency. Hardware manufacturers, such as NVIDIA, are now offering network switches that deliver several-fold improvements in QoS compared to previous-generation hardware. For intra-machine networking, NVSwitch [nvlink-nvswitch] provides a bandwidth of up to 900 Gb/s, surpassing PCIe’s bandwidth of only up to tens of Gb/s combined with high latency [li2019evaluating]. Furthermore, for Infiniband-based intra-rack networking, NVIDIA Quantum switches now offer up to 400 Gb/s bandwidth per port with ultra-low latency using GPU direct RDMA. In contrast, standard Infiniband offers only up to 50 Gb/s bandwidth while experiencing relatively high latency (up to tens of microseconds) [Infiniband-eval]. Finally, NVIDIA Spectrum Ethernet switches provide up to 800 Gb/s bandwidth for inter-rack networking. Note that ethernet switches suffer from high latency. However, current schedulers lack awareness of such high-speed network hardware, which could be harnessed for improved performance. This underscores the demand for a novel approach to DL cluster scheduling.
III-D Limitations of Current SOTA DL Cluster Schedulers
@@
Current SOTA schedulers like Tiresias [tiresias] and Gandiva [gandiva] have several shortcomings such as: (i) leveraging only limited knowledge of the datacenter network: due to the lack of awareness of the various network tiers, e.g., the medium-bandwidth GPU direct RDMA that exists within a rack, these schedulers often end up making a sub-optimal job placement decision, in this case, mapping the job across multiple racks; (ii) a strict consolidation policy: restricting the job placement to either the fewest machines possible (for high skewed models) may lead to high queueing delays in case the former is not available. This necessitates an intermediate approach, allowing for a gradual and graceful “relaxation” of consolidation, particularly for moderately sensitive jobs based on queueing delay and network sensitivity. (iii) Sub-optimal job priority: Tiresias employs a generalized version of Least Attained Service (LAS) [las] called “Discretized 2D-LAS” (2DAS) [tiresias] to determine job priority. The 2DAS value for a job is calculated by multiplying the time for which it has been running by the number of GPUs it utilizes. Tiresias prioritizes jobs based on their 2DAS values, with lower values indicating higher priority. Consequently, a job with the highest 2DAS value is least likely to be scheduled for execution and most likely to be preempted out. However, a job with unfavorable network placement and high network sensitivity may have been running for an extended period but may not have made significant progress due to the excessive network latency associated with its placement. This would lead to a high 2DAS value and make it less likely to receive resource offers that are consolidated, as jobs with lower 2DAS values would receive the most favorable offers.
@@
III-E A high-fidelity simulator for DDL scheduling research
@@
The exorbitant cost of deep learning GPUs, with several GPUs required to create a large deep-learning cluster, poses a significant challenge in conducting real-world experiments. This, combined with the need to run such experiments for extended periods, can escalate the cost to the hundreds of thousands of dollars. Even utilizing public cloud resources may prove uneconomical as deep learning machines like the AWS P4 instances incur hourly charges of more than \$32/hour [aws_pricing]. Further complicating matters, research based on cutting-edge network hardware (such as the NVIDIA Quantum and Spectrum switches) necessitates a simulation platform due to its limited availability. To the best of our knowledge, no public cloud provider offers access to such hardware. To address these challenges, research efforts like [rashidi2022themis] have turned to ASTRA-sim as a viable tool for simulating the latest network hardware. However, ASTRA-sim is a single DDL job simulator.
@@
Existing multi-job DDL cluster scheduler simulators, such as [themis-nsdi], suffer from low accuracy due to their unrealistic approach to network latency modeling. For example, [themis-nsdi] implements static slowdown penalties of job placement on the rack and the network irrespective of the DNN model or network hardware. But, to achieve high confidence in the simulation results, a DL cluster scheduler simulator must dynamically determine the accurate network latency of a job based on specific placement within the cluster and the type of network hardware used to accurately reflect real-world conditions. In response to this requirement, we propose a “multi-job” DL cluster simulator that employs ASTRA-sim to accurately determine network latency.
IV DALLY DESIGN
@@
We now describe Dally, a novel network-placement sensitive DL cluster scheduler that leverages the principles of delay scheduling and a novel network-sensitive preemption policy. In the following sections, we delve into the details of our system design, including the simulation methodology employed to evaluate its effectiveness.
IV-A Overview
IV-A1 System goal
@@
The system processes a stream of DL/DDL jobs that need to be scheduled for execution. These jobs exhibit heterogeneity in terms of the DNN model, the number of iterations required to attain the target accuracy, the compute time per iteration, and the GPU demand per job. The system may encounter jobs arriving in batches so that the aggregate GPU demand across all jobs would normally surpass the total number of GPUs available in the cluster. Alternatively, jobs may arrive according to a Poisson distribution. Our goal is to construct a DL cluster scheduler that aligns with the target objectives and assumptions elaborated upon in the subsequent sections. This scheduler will function within a cluster comprising homogeneous GPUs connected via latest-generation network hardware in a hierarchical structure, as discussed in Section III. Generally, the scheduler is intended for an MLaaS provider who may have little to no knowledge about the DNN models being trained including any historical data associated with it.
IV-A2 Objective
@@ IV-A3 Constraints
  3. 3.
    Unknown job submission pattern: The scheduler lacks information about the order and frequency of job arrivals.
  4. 4.
@@
IV-B Design of the Dally Scheduler
@@
As stated earlier, Dally is designed to minimize cloud expenditure. To achieve this objective, it uses delay scheduling for job placement and employs a novel preemption strategy that incorporates network sensitivity. When a job is submitted to the cluster, the global scheduler (refer to Fig. 3) initially places it into a wait queue. The scheduler then periodically generates “resource offers” to jobs in the wait queue based on the availability of cluster resources. The job’s local scheduler (Fig. 3) makes a decision to accept or reject the offer based on delay timers which is maintained for each network dimension. Upon acceptance, the job is moved to the run queue, where it executes until it is either preempted or completes its execution. Upon preemption, the job is reintroduced into the wait queue (after saving its current state) and remains there until it accepts another resource offer. The job then executes from its last saved state which consists of the model parameters, the state of the optimizer (e.g., current step size) and the number of iterations completed. Upon restart, the model state is first restored from the persistent storage and the job executes from the last iteration completed. Note that the restore time is negligible compared to the long running times of DL jobs. An instance of Dally’s mechanisms governing preemption and job placement is elaborated upon in the following sections.
Figure 3: Scheduling scheme.
IV-B1 Preemption Priority